# Where Visual Speech Meets Language: VSP-LLM Framework for Efficient and Context-Aware Visual Speech Processing

### v1

@@ Abstract
In visual speech processing, context modeling capability is one of the most important requirements due to the ambiguous nature of lip movements. For example, homophenes, words that share identical lip movements but produce different sounds, can be distinguished by considering the context. In this paper, we propose a novel framework, namely Visual Speech Processing incorporated with LLMs (VSP-LLM), to maximize the context modeling ability by bringing the overwhelming power of LLMs. Specifically, VSP-LLM is designed to perform multi-tasks of visual speech recognition and translation, where the given instructions control the type of task. The input video is mapped to the input latent space of a LLM by employing a self-supervised visual speech model. Focused on the fact that there is redundant information in input frames, we propose a novel deduplication method that reduces the embedded visual features by employing visual speech units. Through the proposed deduplication and Low Rank Adaptors (LoRA), VSP-LLM can be trained in a computationally efficient manner. In the translation dataset, the MuAViC benchmark, we demonstrate that VSP-LLM can more effectively recognize and translate lip movements with just 15 hours of labeled data, compared to the recent translation model Anwar et al. (2023) trained with 433 hours of labeld data. The code and demo are available at: https://github.com/Sally-SH/VSP-LLM
footnotetext: $*{}^{*}$qual Contribution. ${}^{\dagger}$orresponding Author.
@@ 1 Introduction
Along with audio, visual speech (e.g., lip movements) plays a critical role in human communication. With the increasing acknowledgment of the importance of visual speech, a diverse range of visual-based speech processing technologies Assael et al. (2016); Petridis and Pantic (2016); Chung and Zisserman (2017a); Ma et al. (2021a, 2022b) is emerging. For instance, Visual Speech Recognition (VSR) Kim et al. (2021); Ma et al. (2022a); Yeo et al. (2023a) allows for the identification of spoken words through the observation of lip movements alone, without the need for audio access. Most recently, the exploration has begun into Visual Speech Translation (VST) Cheng et al. (2023), which directly generates translated text in the target language from the input lip movements of the source language.
@@
To this end, in this paper, we propose a new framework named Visual Speech Processing incorporated with LLM (VSP-LLM) that learns the seamless embedding of visual speech into the learned text space of LLMs. VSP-LLM employs a self-supervised visual speech model to embed the input visual speech into phoneme-level representations, where the derived phonetic information can be effectively associated with text Zhang et al. (2022b). Moreover, to reduce the computational burden in training along with LLMs, we propose a novel deduplication method that reduces the input sequence lengths of LLMs. Concretely, we employ visual speech units, the discretized representations of the features from a self-supervised model, as indicators for overlapped information between sequences. As the visual speech units can be regarded as pseudo-text Lakhotia et al. (2021), the visual speech features assigned to the same visual speech units are averaged to reduce the processing of redundant information and improve computational efficiency. Through analysis, we show that we can reduce the sequence length by about 50% with the proposed deduplication without any performance degradation. Finally, the proposed VSP-LLM is jointly trained to perform VSR and VST with a single model which is the first explored in this paper. We show that by bringing the powerful context modeling ability into visual speech processing, we achieve state-of-the-art performances in both VSR and VST, and comparable performances with the previous methods even if we utilize just 15 hours of supervised data.
The key contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this is the first work to incorporate visual speech modeling with LLMs and achieve state-of-the-art performances in VSR and VST. 2) This is the first to work to develop a unified visual speech processing model that can perform both VSR and VST with a single trained model. 3) We propose a novel visual speech deduplication that significantly improves computational efficiency. 4) We show that the proposed VSP-LLM can perform multi-tasks with superior performances even in limited training resource situations, just with 15 hours of labeled data by outperforming the recent translation model Anwar et al. (2023).
@@ 2.1 Visual Speech Processing
Visual speech processing technologies are mainly comprised of two parts, VSR and VST. VSR is a task to recognize the language content by watching lip movements, without any sound. The VSR technologies have greatly progressed with the development of deep learning. Early works Chung and Zisserman (2017b); Stafylakis and Tzimiropoulos (2017); Petridis et al. (2017, 2018) utilize the CNN He et al. (2016) and the RNN Chung et al. (2014); Hochreiter and Schmidhuber (1997) to devise a word-level VSR system. To expand the VSR systems into sentence-level, Chung et al. (2017); Afouras et al. (2018) have utilized a multi-stage pipeline to automatically collect large-scale VSR data. Based on the large-scale VSR datasets, researchers Serdyuk et al. (2022); Ma et al. (2021b) have developed the VSR systems from the perspective of architecture, especially the Transformer Vaswani et al. (2017) have greatly improved the performance of VSR by enabling to capture of the context between any two positions of lip sequences. Moreover, the multimodal learning strategies Zhao et al. (2020); Afouras et al. (2020); Ren et al. (2021); Ma et al. (2021a); Kim et al. (2021, 2022); Yeo et al. (2023b) have attempted to complement the insufficient visual speech representations by utilizing audio information. A recent self-supervised model known as AV-HuBERT Shi et al. (2022), has significantly improved the visual speech representations by predicting the pseudo-label assigned from clustering audio-visual features, with a mask-prediction task like BERT Devlin et al. (2019). According to the advancement of the VSR system, now we can recognize lip movements quite accurately through state-of-the-art VSR models such as AV-HuBERT. Building upon this, the exploration for VST has begun by introducing a Multilingual Audio-Visual Corpus (MuAViC) Anwar et al. (2023) dataset and constructing a VST Cheng et al. (2023).
Despite these research efforts, the development of visual speech processing systems enabling multi-task via a unified model, such as VSR and VST, has been never explored in the previous visual speech processing literature. Hence, the objective of this paper is to develop a unified model to perform multi-tasks, including VSR and VST, by utilizing a rich context modeling ability of LLMs.
Figure 1: Illustration of our VSP-LLM framework. Visual speech representations encoded from the visual encoder are mapped to visual speech units. Then the visual speech representations are reduced through averaging based on the mapped visual speech units. These reduced representations are fed into the Large Language Model (LLM) along with text instructions. Throughout this process, the LLM is kept frozen, with only a small number of QLoRA parameters being fine-tuned.
@@ 2.2 Integration of speech models and LLMs
LLMs have shown remarkable success in various tasks due to their extensive linguistic knowledge and contextual understanding. While leveraging such inherent advantages of LLMs, several studies have tried to seamlessly integrate text-based knowledge with other modalities, particularly in the audio speech domain. For example, AudioPaLM Rubenstein et al. (2023) has been proposed to build a unified model interacting between text language and audio speech. To naturally bridge the gap between the two modalities, AudioPaLM has developed a multimodal vocabulary composed of discrete tokens representing both text and speech. Fathullah et al. Fathullah et al. (2023) have employed LLaMA as a speech recognition decoder, so that the speech sequence features obtained from a conformer encoder were designed to be directly mapped into text tokens, the domain of LLaMA. Moreover, Wu et al. Wu et al. (2023a) have tried to address the inherent problem of mismatched sequence lengths between speech signals and text, while taking LLaMA as a speech translation decoder. So, they have compressed the speech sequence feature and matched its sequence length with that of the text.
However, while the existing studies have primarily focused on incorporating LLMs with the audio speech modality, the exploration of such integration for visual speech processing remains unexplored. In this paper, we propose a novel framework that integrates visual speech processing with LLMs. Specifically, we attempt to mitigate the homophenes problem, one of the key challenges in the field of visual speech processing, by leveraging the rich context modeling capabilities of LLMs. Additionally, to address the training load issues arising from the integration of the visual speech model and the Language Model (LLM), we introduce the concept of a visual speech unit. Through the implementation of visual speech units, we propose a novel visual speech deduplication method that compresses redundant representations while preserving contextual information.
@@ 3.3 Multi-task Learning with Instruction
where $X$ is input video, $I$ is instruction used, $y^{<l}$ is the previous predictions, and $L$ is the length of ground truth. Please note that this is the first work exploring a unified framework of VSR and VST. For training, we employ the recently proposed QLoRA Dettmers et al. (2023) to further relieve the computational load in training LLM.
@@ 4.1 Dataset
Lip Reading Sentences 3 (LRS3) Afouras et al. (2018) is the most widely-used dataset for VSR, which comprises 433 hours of English audio-visual speech corpus with transcription data. These corpora are collected from the TED and TEDx talks. We utilize the LRS3 dataset to and measure the VSR performance of the proposed unified model. Multilingual Audio-Visual Corpus (MuAViC) Anwar et al. (2023) is a multilingual audio-visual dataset designed for speech recognition and speech-to-text translation. It includes 1200 hours of audio-visual corpus in 9 languages, providing full transcriptions and covering 6 English-to-X translations, as well as 6 X-to-English translation directions. To evaluate the VST performance of our model, we utilize English-to-X translation data from MuAViC dataset, where X can be among four languages, Spanish (Es), French (Fr), Portuguese (Pt), and Italian (It). For training our model, we combine LRS3 Afouras et al. (2018) and English-to-X translation data of MuAViC Anwar et al. (2023).
@@
1,759 433 26.9
Table 1: The performance comparisons with state-of-the-art supervised and self-supvervised VSR methods. Compared to the self-supervised methods, the proposed unified model, which can perform both VSR and VST, achieves state-of-the-art recognition performances. Furthermore, our method obtains comparable performances even with much less amount of training data. Please note that the only method which can perform recogntion and translation at once is the proposed model.
  Method Modality Labeled data(hrs) BLEU normal-$\uparrow$
@@
Anwar et al. (2023) AV 433 20.7 25.3 20.5 26.6 23.3
AV-HuBERT V 433 21.5 26.5 26.7 24.5 24.8
Cascaded (AV-HuBERT + MT) V 433 22.2 23.9 24.4 23.5 23.5
@@
Proposed Method V 433 29.4 33.0 33.7 33.8 32.5
Table 2: The experimental results showing the effectiveness on VST, English to target language (En-X) speech translation tasks, on MuAViC.
@@ 4.2 Implementation Details
Training and evaluation. We follow AV-HuBERT Ren et al. (2021) except for the number of updates and learning rate. We conduct training with 15K updates and a learning rate of $1e^{-3}$ for LRS3 1h, 5h, 10h, and 30K updates with a learning rate of $5e^{-4}$ for LRS3 30h and 433h. Adam optimizer is employed for training with $\beta_{1}=0.9$ and $\beta_{2}=0.98$, utilizing a tri-stage learning rate scheduler. The training process is executed on 8 3090 RTX GPUs. For decoding, we use a beam search with a beam width of 20 and a length penalty of 0. We assess the performance of our model using Word Error Rate (WER) for the VSR task and BLEU score Papineni et al. (2002) for the VST task. We use total FLOPs per epoch as a metric to measure the model operation count during training.
@@ 4.3.1 Comparison with State-of-the-arts
In this subsection, we compare the proposed unified model with state-of-the-art VSR methods and VST methods. Please note that the proposed model can perform multi-tasks VSR and VST with a single trained model while the other models need a single model per specific task.
Table 1 presents the performance comparisons of the proposed method with state-of-the-art VSR methods on the LRS3 dataset. The top section of Table 1 outlines the performance of current supervised approaches that depend on extensive labeled training data, while the lower section presents a comparison with other self-supervised methods. Table 1 demonstrates that our approach achieves performance on par with others by employing just 30 hours of labeled data, despite the proposed unified model’s ability to handle multiple tasks—VSR and VST—simultaneously. Additionally, with the use of 433 hours of data, our method reaches state-of-the-art performance, achieving a WER of 26.9%, surpassing other self-supervised approaches. Moreover, Table 1’s upper part shows that the existing supervised methods record exceptional performance using (tens of) thousands of labeled data. However, it is important to highlight that the proposed unified model can obtain comparable performances to several supervised methods.
Table 2 presents the comparison results of VST performance. We construct two baseline models for comparison. The first, AV-HuBERT, is trained similarly to our approach, utilizing both VSR and VST datasets. The second model is a cascaded system that incorporates a pre-trained AV-HuBERT for VSR with a neural machine translation model Fan et al. (2021). Through this comparison, our proposed unified model demonstrates superior VST performance across four English-to-X translation tasks, achieving BLEU scores of 29.4, 33.0, 33.7, and 33.8 for English to Italian, French, Portuguese, and Spanish, respectively. Moreover, it is worth noting that the proposed method achieves a 23.5 BLEU score on average with only 15 hours of labeled data, outperforming the audio-visual speech translation model Anwar et al. (2023) trained with 433 hours of labeled data.
Figure 3: The qualitative results showing that the contextual modeling ability of LLM, which is adopted in our method, can improve the homophene problem and other confusing cases. The red and blue words indicate the wrong predictions from AV-HuBERT. However, as shown in the examples, the proposed method can generate correct words by considering the entire context (e.g.,barrow’ to ‘marrow’).

### v2

@@ Abstract
In visual speech processing, context modeling capability is one of the most important requirements due to the ambiguous nature of lip movements. For example, homophenes, words that share identical lip movements but produce different sounds, can be distinguished by considering the context. In this paper, we propose a novel framework, namely Visual Speech Processing incorporated with LLMs (VSP-LLM), to maximize the context modeling ability by bringing the overwhelming power of LLMs. Specifically, VSP-LLM is designed to perform multi-tasks of visual speech recognition and translation, where the given instructions control the type of task. The input video is mapped to the input latent space of an LLM by employing a self-supervised visual speech model. Focused on the fact that there is redundant information in input frames, we propose a novel deduplication method that reduces the embedded visual features by employing visual speech units. Through the proposed deduplication and Low Rank Adaptation (LoRA), VSP-LLM can be trained in a computationally efficient manner. In the translation dataset, the MuAViC benchmark, we demonstrate that VSP-LLM trained on just 30 hours of labeled data can more effectively translate lip movements compared to the recent model trained with 433 hours of data.
footnotetext: Equal Contribution. Corresponding Author.
@@ 1 Introduction
Along with audio, visual speech (e.g., lip movements) plays a critical role in human communication. With the increasing acknowledgment of the importance of visual speech, a diverse range of visual-based speech processing technologies Assael et al. (2016); Petridis and Pantic (2016); Chung and Zisserman (2017a); Ma et al. (2021a, 2022b); Yemini et al. (2024) is emerging. For instance, Visual Speech Recognition (VSR) Kim et al. (2021); Ma et al. (2022a); Yeo et al. (2023a) allows for the identification of spoken words through the observation of lip movements alone, without the need for audio access. Most recently, the exploration has begun into Visual Speech Translation (VST) Cheng et al. (2023), which directly generates translated text in the target language from the input lip movements of the source language.
@@
To this end, in this paper, we propose a new framework named Visual Speech Processing incorporated with LLM (VSP-LLM) that learns the seamless embedding of visual speech into the learned text space of LLMs. VSP-LLM employs a self-supervised visual speech model to embed the input visual speech into phoneme-level representations, where the derived phonetic information can be effectively associated with text Zhang et al. (2022b). Moreover, to reduce the computational burden in training along with LLMs, we propose a novel deduplication method that reduces the input sequence lengths of LLMs. Concretely, we employ visual speech units, the discretized representations of the features from a self-supervised model, as indicators for overlapped information between sequences. As the visual speech units can be regarded as pseudo-text Lakhotia et al. (2021), the visual speech features assigned to the same visual speech units are averaged to reduce the processing of redundant information and improve computational efficiency. Through our analysis, we show that the sequence length can be reduced by approximately 50% using the proposed deduplication, with minimal performance degradation. Finally, the proposed VSP-LLM is jointly trained to perform VSR and VST with a single model which is the first explored in this paper. We show that by bringing the powerful context modeling ability into visual speech processing, we achieve state-of-the-art performances in both VSR and VST when using the LRS3 Afouras et al. (2018) and MuAViC Anwar et al. (2023) datasets as training data. Additionally, our VSP-LLM trained with just 30 hours of data outperforms the recent translation model used 433 hours of training data.
The key contributions of this paper can be summarized as follows: 1) To the best of our knowledge, this is the first work to incorporate visual speech modeling with LLMs and achieve state-of-the-art performances in VSR and VST. 2) This is the first to work to develop a unified visual speech processing model that can perform both VSR and VST with a single trained model. 3) We propose a novel visual speech deduplication that significantly improves computational efficiency. 4) We show that the proposed VSP-LLM can perform multi-tasks with superior performances even in limited training resource situations, just with 30 hours of labeled data by outperforming the recent translation model.
@@ 2.1 Visual Speech Processing
Visual speech processing technologies are mainly comprised of two parts, VSR and VST. VSR is a task to recognize the language content by watching lip movements, without any sound. The VSR technologies have greatly progressed with the development of deep learning. Early works Chung and Zisserman (2017b); Stafylakis and Tzimiropoulos (2017); Petridis et al. (2017, 2018) utilize the CNN He et al. (2016) and the RNN Chung et al. (2014); Hochreiter and Schmidhuber (1997) to devise a word-level VSR system. To expand the VSR systems into sentence-level, Chung et al. (2017); Afouras et al. (2018) have utilized a multi-stage pipeline to automatically collect large-scale VSR data. Based on the large-scale VSR datasets, researchers Serdyuk et al. (2022); Ma et al. (2021b) have developed the VSR systems from the perspective of architecture, especially the Transformer Vaswani et al. (2017) have greatly improved the performance of VSR by enabling to capture of the context between any two positions of lip sequences. Moreover, the multimodal learning strategies Zhao et al. (2020); Afouras et al. (2020); Ren et al. (2021); Ma et al. (2021a); Kim et al. (2021, 2022); Yeo et al. (2023b) have attempted to complement the insufficient visual speech representations by utilizing audio information. A recent self-supervised model known as AV-HuBERT Shi et al. (2022), has significantly improved the visual speech representations by predicting the pseudo-label assigned from clustering audio-visual features, with a mask-prediction task like BERT Devlin et al. (2019). According to the advancement of the VSR system, we can now recognize lip movements quite accurately through state-of-the-art VSR models such as AV-HuBERT. Building upon this, the exploration for VST has begun by introducing a Multilingual Audio-Visual Corpus (MuAViC) Anwar et al. (2023) dataset and constructing a VST Cheng et al. (2023).
Despite these research efforts, the development of visual speech processing systems enabling multi-task via a unified model, such as VSR and VST, has never been explored in the previous visual speech processing literature. Hence, the objective of this paper is to develop a unified model to perform multi-tasks, including VSR and VST, by utilizing a rich context modeling ability of LLMs.
Figure 1: Illustration of our VSP-LLM framework. Visual speech representations encoded from the visual encoder are mapped to visual speech units. Then the visual speech representations are reduced through averaging based on the mapped visual speech units. These reduced representations are fed into the LLM along with text instructions.
@@ 2.2 Integration of speech models and LLMs
LLMs have shown remarkable success in various tasks due to their extensive linguistic knowledge and contextual understanding. While leveraging such inherent advantages of LLMs, several studies have tried to seamlessly integrate text-based knowledge with other modalities, particularly in the audio speech domain. For example, AudioPaLM Rubenstein et al. (2023) has been proposed to build a unified model interacting between text language and audio speech. To naturally bridge the gap between the two modalities, AudioPaLM has developed a multimodal vocabulary composed of discrete tokens representing both text and speech. Fathullah et al. Fathullah et al. (2023) have employed LLaMA as a speech recognition decoder so that the speech sequence features obtained from a conformer encoder were designed to be directly mapped into text tokens, the domain of LLaMA. Moreover, Wu et al. Wu et al. (2023a) have tried to address the inherent problem of mismatched sequence lengths between speech signals and text, while taking LLaMA as a speech translation decoder. So, they have compressed the speech sequence feature and matched its sequence length with that of the text.
However, while the existing studies have primarily focused on incorporating LLMs with the audio speech modality, the exploration of such integration for visual speech processing remains unexplored. In this paper, we propose a novel framework that integrates visual speech processing with LLM. Specifically, we attempt to mitigate the homophenes problem, one of the key challenges in the field of visual speech processing, by leveraging the rich context modeling capabilities of LLM. Additionally, to address the training load issues arising from the integration of the visual speech model and LLM, we introduce the concept of a visual speech unit. Through the implementation of visual speech units, we propose a novel visual speech deduplication method that compresses redundant representations while preserving contextual information.
@@ 3.3 Multi-task Learning with Instruction
where $X$ is input video, $I$ is instruction used, $y^{l}$ is the $l$-th text token of the ground truth sentence, $y^{<l}$ is the previous predictions, and $L$ is the length of ground truth. Please note that this is the first work exploring a unified framework of VSR and VST. For training, we employ the recently proposed QLoRA Dettmers et al. (2023) to further relieve the computational load in training LLM.
@@ 4.1 Dataset
Lip Reading Sentences 3 (LRS3) Afouras et al. (2018) is the most widely-used dataset for VSR, which comprises 433 hours of English audio-visual speech corpus with transcription data. These corpora are collected from the TED and TEDx talks. We utilize the LRS3 dataset to measure the VSR performance of the proposed unified model.
Multilingual Audio-Visual Corpus (MuAViC) Anwar et al. (2023) is a multilingual audio-visual dataset designed for speech recognition and speech-to-text translation. It includes 1200 hours of audio-visual corpus in 9 languages, providing full transcriptions and covering 6 English-to-X translations, as well as 6 X-to-English translation directions. To evaluate the VST performance of our model, we utilize English-to-X translation data from MuAViC dataset, where X can be among four languages, Spanish (Es), French (Fr), Portuguese (Pt), and Italian (It). For training our model, we combine the LRS3 dataset and English-to-X translation data of MuAViC.
@@
VSP-LLM(FT) 1,759 433 25.4
Table 1: The performance comparisons with state-of-the-art VSR methods. Compared to the self-supervised methods, the proposed VSP-LLM, which can perform both VSR and VST, achieves state-of-the-art recognition performances. We also evaluate the performance of a fine-tuned VSP-LLM(FT) with an unfrozen visual encoder.
  Method Labeled data(hrs) BLEU $\uparrow$
@@
Anwar et al. (2023) 433 15.1 16.8 15.1 19.2 16.6
AV-HuBERT 433 16.6 19.4 17.4 21.7 18.8
Cascaded (AV-HuBERT + MT) 433 17.6 19.5 17.4 22.4 19.2
@@
VSP-LLM 433 17.9 22.3 18.7 22.7 20.4
VSP-LLM(FT) 433 17.7 22.2 19.4 22.4 20.4
Table 2: Experimental results for English to target language (En-X) translation on the MuAViC benchmark.
@@ 4.2 Implementation Details
Training and evaluation. We follow AV-HuBERT Ren et al. (2021) except for the number of updates and learning rate. We conduct training with a learning rate of $5e^{-4}$ and the number of updates is 15K updates for LRS3 1h, 5h, 10h, and 30K updates for LRS3 30h and 433h. For VSP-LLM (FT), the visual encoder is frozen for the first 18K steps and then unfrozen afterward. Adam optimizer is employed for training with $\beta_{1}=0.9$ and $\beta_{2}=0.98$, utilizing a tri-stage learning rate scheduler. The training process is executed on 8 3090 RTX GPUs. For decoding, we use a beam search with a beam width of 20 and a length penalty of 0. We assess the performance of our model using Word Error Rate (WER) for the VSR task and BLEU score Papineni et al. (2002) for the VST task. We use total FLOPs per epoch as a metric to measure the model operation count during training.
@@ 4.3.1 Comparison with State-of-the-arts
In this subsection, we compare the proposed unified model with state-of-the-art VSR and VST methods. Please note that the proposed model can perform multi-tasks VSR and VST with a single trained model while the other models need a single model per specific task.
Table 1 presents the performance comparisons of the proposed method with state-of-the-art VSR methods on the LRS3 dataset. The top section of Table 1 outlines the performance of current supervised approaches that depend on extensive labeled training data, while the lower section presents a comparison with other self-supervised methods. Table 1 demonstrates that our approach achieves performance on par with others by employing just 30 hours of labeled data, despite the proposed unified model’s ability to handle multiple tasks—VSR and VST—simultaneously. When employing 433 hours of training data, our method achieves a WER of 26.7%. By fine-tuning the VSP-LLM(FT) with an unfrozen visual encoder, we further enhance our performance, achieving a WER of 25.4%, surpassing other self-supervised approaches. Moreover, Table 1’s upper part shows that the existing supervised methods record exceptional performance using (tens of) thousands of labeled data. However, it is important to highlight that the proposed unified model can obtain comparable performances to several supervised methods.
Table 2 presents the comparison results of VST performance. We construct two baseline models for comparison. The first, AV-HuBERT, is trained similarly to our approach, utilizing both VSR and VST datasets. The second model is a cascaded system that incorporates a pre-trained AV-HuBERT for VSR with a neural machine translation model Fan et al. (2021). Through this comparison, our proposed VSP-LLM demonstrates superior VST performance across four English-to-X translation tasks, achieving BLEU scores of 17.9, 22.3, 18.7, and 22.7 for English to Italian, French, Portuguese, and Spanish, respectively. The VSP-LLM(FT) shows a better performance 19.4 BLUE score on translation from English to Portuguese and comparable performances in other languages. Moreover, it is worth noting that the proposed method achieves an 18.2 BLEU score on average with only 30 hours of labeled data, outperforming the bilingual speech translation model Anwar et al. (2023) trained with 433 hours of labeled data.
Figure 3: The qualitative results showing that the contextual modeling ability of LLM, which is adopted in our method, can improve the homophene problem and other confusing cases. The red and blue words indicate the wrong predictions from AV-HuBERT. However, as shown in the examples, the proposed method can generate correct words by considering the entire context (e.g.,i’ to ‘eye’).